OpenStudy (anonymous):

one-eighth can be written as 2 to the what power?

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\] so \[2^{-3}=\frac{1}{2^3}=\frac{1}{8}\]

OpenStudy (anonymous):

so... \[2^-3\]

OpenStudy (anonymous):

question is "what power" answer is "-3"

OpenStudy (anonymous):

thanks